# RM-RL: Role-Model Reinforcement Learning for Precise Robot Manipulation

### v1

@@ III-D Cell Plate Localization
Then, the yaw angle $\psi$ can be estimated from $\mathcal{M}$ by applying Principal Component Analysis (PCA) [52], where the first principal component indicates the major directions. Finally, we use the combination of $\mathbf{p}_{world}$ and $\psi$ as the pick pose of the end-effector in the robotic arm to finish the picking task.
@@ IV Experiments
IV-A Hardware
For the hardware, as illustrated in Fig. 4, we use a 6-DOF Ufactory X-ARM 6 with a bio gripper as our platform to finish the designed experiments. RealSense D435 is mounted overhead to provide depth and RGB visual perception, enabling the system to detect and localize the cell plate in the workspace. Two target positions, marked as red stars, are defined in the setup: one is located on the table surface (where the cell plate is placed), and the other is on the shelf slot. This configuration enables evaluation of reinforcement learning–based manipulation policies under real-world conditions, requiring accurate transfers between different spatial levels. For the calibration of the overhead RealSense camera, we first calibrate the camera on the robotic wrist through chessboard calibration method [53] to get the transformation matrix $T^{wrist}_{world}$. Then, the wrist camera and the overhead camera localize the chessboard simultaneously, to get the transformation matrix $T_{wrist}^{camera}$. Finally, we get the transformation matrix $T^{camera}_{world}=T_{wrist}^{camera}\cdot T^{wrist}_{world}$.
IV-B Baseline
To verify the effectiveness of the proposed method, we compare the performance of the standard RL [22], standard RL with replay buffer (RL + Replay Buffer) [12], the proposed Role-Model Reinforcement Learning (RM-RL), and pretrained RM-RL. To ensure fair comparisons, all RL methods use the same policy network and are trained with the same policy gradient-based method and training configurations. Since the selected task is relatively simple, we adopt policy gradient for optimization instead of more advanced algorithms such as PPO or Actor-Critic. It is worth noting that our proposed role-model strategy is orthogonal to these algorithms and can be readily integrated into them in future work. For the pretrained RL + MR, we collect about 200 samples, which are labeled through the proposed role-model strategy, and the policy model is trained before the beginning of real-world RL training. To ensure consistency, all policy networks were trained and evaluated on the same PC, equipped with an NVIDIA RTX 4060 GPU (6 GB).
IV-C Training Performance
@@
Methods $T_{\tau}^{10}$ Average Reward
Standard RL [22] 45 0.948 (0.029)
RL + Replay Buffer [12] 36 0.953 (0.031)
RM-RL 35 0.956 (0.027)
Pretrained RM-RL 30 0.969 (0.023)
TABLE I: Quantitative experimental results of different RL methods evaluated by $T_{\tau}^{10}$ and average reward.
For quantitative analysis, we use two metrics to evaluate the training performance. Similar to the metric $J^{eff.}$ [21], we record the training step index at which the learned policy achieves a predefined reward threshold for the tenth occurrence. The calculation of the metric is shown as follows:
@@
Table I summarizes the quantitative experimental results. In particular, Pretrained RM-RL achieves the best results, requiring only 30 steps to reach the threshold for the 10th time (a 33% reduction from Standard RL) and attaining the highest average reward of 0.969 with the lowest variance, indicating faster convergence and more stable learning. Although the RL with Relay Buffer demonstrates competitive performance, it exhibits the largest standard deviation in experimental results, indicating instability in training.
IV-D Real-World Performance
@@
To evaluate the pick and place performance in the real-world environment, we conduct two experiments. The first experiment follows the training process, putting one cell plate into the given box of a similar size. We provide 10 random initial positions of the cell plate. The robot needs to pick the cell plate from the initial position and place it in the target position. We record the average reward, translation error $e_{trans}$, and rotation error $e_{rot}$ in Table II of the compared methods for evaluation. From the experimental results, the Pretrained RM-RL achieves the highest average reward (0.978). Its translational error is reduced to 2.1 mm, representing a 42% improvement over Standard RL (3.6 mm) and a 64% reduction compared with RL + Replay Buffer (5.8 mm). For rotational accuracy, Pretrained RM-RL (0.60°) is nearly identical to Standard RL (0.58°) and 41% better than RL + Replay (1.01°). Overall, these results demonstrate that pretraining using role-model-labeled data significantly enhances real-world pick-and-place performance.
@@
Methods Average Reward $e_{trans}$/$m$ $e_{rot}$$\mathrm{\SIUnitSymbolDegree}$
Standard RL 0.964 (0.024) 3.6 (0.27) 0.58 (0.49)
RL + Replay Buffer 0.944 (0.030) 5.8 (0.33) 1.01 (0.43)
RM-RL 0.970 (0.025) 3.0 (0.30) 0.89 (0.38)
Pretrained RM-RL 0.978 (0.012) 2.1 (0.12) 0.60 (0.45)
@@
TABLE II: Performance comparison of different reinforcement learning strategies in real-world robotic picking and placing tasks, evaluated by average reward, position error, and rotation error.
@@
The other experiment is placing a cell plate onto the cell plate shelf, given random initial states. The aim of this experiment is to evaluate whether the predicted pose adjustment is also effective for other tasks that also need precise operations, and the influence of the small errors on these RL-based methods. The robotic arm autonomously picks the cell plate with the localization information and the adjustment pose from RL. A fixed trajectory is designed to guarantee the cell plate can be put on the cell plate shelf with a proper grasping pose. We evaluate the performance of different RL-based methods by measuring the success rate over ten trials. The experiment results are summarized in Table III. From the experimental results, the Standard RL achieves a success rate of 50%, while RL + Replay performs worse with only 40%, demonstrating the replay with distribution transit error may fail the precise tasks. RM-RL also shows unsatisfactory results, achieving only a 70% success rate, despite presenting results comparable to Pretrained RL + MR in the previous experiment. The Pretrained RM-RL succeeds in all trials, demonstrating the effectiveness of the proposed receipt with pretraining in enhancing the robot’s capability for precise tasks.
V Summary and Future Work
@@
Methods Successful Rate
Standard RL 50%
RL + Replay Buffer 40%
RM-RL (Ours) 70%
Pretrained RM-RL (Ours) 100%
TABLE III: Success rates of different reinforcement learning strategies in real-world picking a cell plate on a shelf.
@@
VI Acknowledgments
This work is supported by a Start-up Grant from Nanyang Technological University and jointly funded by the Singapore Ministry of Education (MOE) under a Tier-1 research grant.

### v2

@@ 3.4 Cell Plate Localization
Then, the yaw angle $\psi$ can be estimated from $\mathcal{M}$ by applying Principal Component Analysis (PCA) Abdi and Williams (2010), where the first principal component indicates the major directions. Finally, we use the combination of $\mathbf{p}_{world}$ and $\psi$ as the pick pose of the end-effector in the robotic arm to finish the picking task.
@@ 4 Experiments
4.1 Hardware
For the hardware, as illustrated in Fig. 4, we use a 6-DOF Ufactory X-ARM 6 with a bio gripper as our platform to finish the designed experiments. RealSense D435 is mounted overhead to provide depth and RGB visual perception, enabling the system to detect and localize the cell plate in the workspace. Two target positions, marked as red stars, are defined in the setup: one is located on the table surface (where the cell plate is placed), and the other is on the shelf slot. This configuration enables evaluation of reinforcement learning–based manipulation policies under real-world conditions, requiring accurate transfers between different spatial levels. For the calibration of the overhead RealSense camera, we first calibrate the camera on the robotic wrist through chessboard calibration method Zhang (2002) to get the transformation matrix $T^{wrist}_{world}$. Then, the wrist camera and the overhead camera localize the chessboard simultaneously, to get the transformation matrix $T_{wrist}^{camera}$. Finally, we get the transformation matrix $T^{camera}_{world}=T_{wrist}^{camera}\cdot T^{wrist}_{world}$.
4.2 Baseline
To verify the effectiveness of the proposed method, we compare the performance of the standard RL Williams (1992), standard RL with replay buffer (RL + Replay Buffer) Mnih et al. (2015), the proposed Role-Model Reinforcement Learning (RM-RL), and pretrained RM-RL. To ensure fair comparisons, all RL methods use the same policy network and are trained with the same policy gradient-based method and training configurations. Since the selected task is relatively simple, we adopt policy gradient for optimization instead of more advanced algorithms such as PPO or Actor-Critic. It is worth noting that our proposed role-model strategy is orthogonal to these algorithms and can be readily integrated into them in future work. For the pretrained RL + MR, we collect about 200 samples, which are labeled through the proposed role-model strategy, and the policy model is trained before the beginning of real-world RL training. To ensure consistency, all policy networks were trained and evaluated on the same PC, equipped with an NVIDIA RTX 4060 GPU (6 GB).
4.3 Training Performance
@@
Methods $T_{\tau}^{10}$ Average Reward Standard RL Williams (1992) 45 0.948 (0.029) RL + Replay Buffer Mnih et al. (2015) 36 0.953 (0.031) RM-RL 35 0.956 (0.027) Pretrained RM-RL 30 0.969 (0.023)
Table 1: Quantitative experimental results of different RL methods evaluated by $T_{\tau}^{10}$ and average reward.
For quantitative analysis, we use two metrics to evaluate the training performance. Similar to the metric $J^{eff.}$ Dulac-Arnold et al. (2019), we record the training step index at which the learned policy achieves a predefined reward threshold for the tenth occurrence. The calculation of the metric is shown as follows:
@@
Table 1 summarizes the quantitative experimental results. In particular, Pretrained RM-RL achieves the best results, requiring only 30 steps to reach the threshold for the 10th time (a 33% reduction from Standard RL) and attaining the highest average reward of 0.969 with the lowest variance, indicating faster convergence and more stable learning. Although the RL with Relay Buffer demonstrates competitive performance, it exhibits the largest standard deviation in experimental results, indicating instability in training.
4.4 Real-World Performance
@@
To evaluate the pick and place performance in the real-world environment, we conduct two experiments. The first experiment follows the training process, putting one cell plate into the given box of a similar size. We provide 10 random initial positions of the cell plate. The robot needs to pick the cell plate from the initial position and place it in the target position. We record the average reward, translation error $e_{trans}$, and rotation error $e_{rot}$ in Table 2 of the compared methods for evaluation. From the experimental results, the Pretrained RM-RL achieves the highest average reward (0.978). Its translational error is reduced to 2.1 mm, representing a 42% improvement over Standard RL (3.6 mm) and a 64% reduction compared with RL + Replay Buffer (5.8 mm). For rotational accuracy, Pretrained RM-RL (0.60°) is nearly identical to Standard RL (0.58°) and 41% better than RL + Replay (1.01°). Overall, these results demonstrate that pretraining using role-model-labeled data significantly enhances real-world pick-and-place performance.
@@
Methods Average Reward $e_{trans}$/$m$ $e_{rot}$$\mathrm{\SIUnitSymbolDegree}$ Standard RL 0.964 (0.024) 3.6 (0.27) 0.58 (0.49) RL + Replay Buffer 0.944 (0.030) 5.8 (0.33) 1.01 (0.43) RM-RL 0.970 (0.025) 3.0 (0.30) 0.89 (0.38) Pretrained RM-RL 0.978 (0.012) 2.1 (0.12) 0.60 (0.45)
@@
Table 2: Performance comparison of different reinforcement learning strategies in real-world robotic picking and placing tasks, evaluated by average reward, position error, and rotation error.
@@
The other experiment is placing a cell plate onto the cell plate shelf, given random initial states. The aim of this experiment is to evaluate whether the predicted pose adjustment is also effective for other tasks that also need precise operations, and the influence of the small errors on these RL-based methods. The robotic arm autonomously picks the cell plate with the localization information and the adjustment pose from RL. A fixed trajectory is designed to guarantee the cell plate can be put on the cell plate shelf with a proper grasping pose. We evaluate the performance of different RL-based methods by measuring the success rate over ten trials. The experiment results are summarized in Table 3. From the experimental results, the Standard RL achieves a success rate of 50%, while RL + Replay performs worse with only 40%, demonstrating the replay with distribution transit error may fail the precise tasks. RM-RL also shows unsatisfactory results, achieving only a 70% success rate, despite presenting results comparable to Pretrained RL + MR in the previous experiment. The Pretrained RM-RL succeeds in all trials, demonstrating the effectiveness of the proposed receipt with pretraining in enhancing the robot’s capability for precise tasks.
5 Summary and Future Work
@@
Methods Successful Rate Standard RL 50% RL + Replay Buffer 40% RM-RL (Ours) 70% Pretrained RM-RL (Ours) 100%
Table 3: Success rates of different reinforcement learning strategies in real-world picking a cell plate on a shelf.
@@
6 Acknowledgments
This work is jointly supported by MOE Singapore Tier 1 Grant RG83/25, RS36/24 and a Start-up Grant from Nanyang Technological University.